OpenStudy (anonymous):

NEWTON METHOD FIND SIX DECIMALS : SINX+X-1=0 WHERE X1=0

OpenStudy (babyslapmafro):

First find the derivative

OpenStudy (babyslapmafro):

Also the equation for newton's method is as follows...

OpenStudy (babyslapmafro):

\[x _{n+1}=x_n-\frac{ f(x_n) }{ f'(x_n) }\]

OpenStudy (anonymous):

YEAH I KNOW THE DERIVATE IS COSX+1

OpenStudy (anonymous):

BUT WHATS NEXT PLUG 0 IN THE EQUATION?

OpenStudy (babyslapmafro):

No, approximately, what x-value allows f(x) to equal 0?

OpenStudy (anonymous):

WELL X1=0

OpenStudy (anonymous):

LETTING THIS BE SO PLUG IT IN THE EQ?

OpenStudy (babyslapmafro):

Ok, yes you can plug in zero and it will yield the same result.

OpenStudy (anonymous):

I HAVE FOR X1= -1/2

OpenStudy (babyslapmafro):

should be positive 1/2

OpenStudy (babyslapmafro):

Be careful when plugging problems like these into your calculator...

OpenStudy (anonymous):

I STILL GET NEGATIVE EXAMPLE 0-SIN0+0-1 DIVIDED BY COS0+1

OpenStudy (babyslapmafro):

My mistake, you are correct.

OpenStudy (babyslapmafro):

Now plug in -1/2 in place of xbase 1 and do so until you find a repeating answer....

OpenStudy (babyslapmafro):

Keep replacing xbase1 with the most recent y-value

OpenStudy (babyslapmafro):

Sorry for not being better help

OpenStudy (anonymous):

SO FOR THE NEXT X2= I PLUG -1/2

OpenStudy (babyslapmafro):

yes, and the value given by x2 is x3

OpenStudy (babyslapmafro):

repeat the calculation until you are given a repeating y-value...it helps to use the table feature on a graphing calculator if you have one.